# Wasserstein distances and divergences of order $p$ by quantum channels

### v2

@@ Abstract.
Bunth was supported by the Momentum Program of the Hungarian Academy of Sciences (grant no. LP2021-15/2021); Pitrik was supported by the “Frontline” Research Excellence Programme of the Hungarian National Research, Development and Innovation Office - NKFIH (grant no. KKP133827) and by the Momentum Program of the Hungarian Academy of Sciences (grant no. LP2021-15/2021); Titkos was supported by the Hungarian National Research, Development and Innovation Office - NKFIH (grant no. K115383) and by the Momentum Program of the Hungarian Academy of Sciences (grant no. LP2021-15/2021).; Virosztek was supported by the Momentum Program of the Hungarian Academy of Sciences (grant no. LP2021-15/2021) and by the ERC Synergy Grant no. 810115.
@@ 1.1. Motivation and main result
In this paper, we propose a non-quadratic generalization of the quantum optimal transport problem relying on quantum channels which was introduced in [16]. In particular, we introduce and study $p$-Wasserstein distances and divergences that arise as (linear combinations of) optimal transport costs with specific cost operators mimicking the classical cost $c(x,y)=\left|\left|x-y\right|\right|_{p}^{p}.$ Finally, we prove triangle inequality for quadratic Wasserstein divergences under the sole assumption that an arbitrary one of the states involved is pure. This is a generalization of our previous result in this direction [5].
@@ 1.2. Basic notions, notation
where $\pi$ runs over all possible couplings of $\mu$ and $\nu.$ We say that $\pi\in\mathrm{Prob}(\mathbb{R}^{K}\times\mathbb{R}^{K})$ is a coupling of $\mu$ and $\nu$ (in notation: $\pi\in\mathcal{C}(\mu,\nu)$) if the marginals of $\pi$ are $\mu$ and $\nu,$ that is, $\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}f(x)\mathrm{d}\pi(x,y)=\int_{\mathbb{R}^{K}}f(x)\mathrm{d}\mu(x)$ and $\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}g(y)\mathrm{d}\pi(x,y)=\int_{\mathbb{R}^{K}}g(y)\mathrm{d}\nu(y)$ for all $f,g\in C_{b}(\mathbb{R}^{K}).$ A consequence of the tightness (that is, sequential compactness in the weak topology) of $\mathcal{C}(\mu,\nu)$ and the lower-semicontinuity of $c$ is that there is a coupling (in other words: transport plan) $\pi_{0}\in\mathcal{C}(\mu,\nu)$ that minimizes (1), see, e.g., [49, Thm. 4.1.].
@@ 2.1. A quantum mechanical optimal transport problem generalizing [16]
$=\prod_{k=1}^{K}\mathrm{tr}_{\mathcal{H}\otimes\mathcal{H}^{*}}\left[\Pi\left(\mathrm{d}E_{k}(y_{k})\otimes\mathrm{d}E_{k}^{T}(x_{k})\right)\right]$
$=\mathrm{tr}_{\left(\mathcal{H}\otimes\mathcal{H}^{*}\right)^{\otimes K}}\left[\Pi^{\otimes K}\left(\mathrm{d}E_{1}(y_{1})\otimes\mathrm{d}E_{1}^{T}(x_{1})\otimes\dots\otimes\mathrm{d}E_{K}(y_{K})\otimes\mathrm{d}E_{K}^{T}(x_{K})\right)\right].$ (5)
@@
Note that $c$ is a non-negative Borel measurable function on $\mathbb{R}^{2K},$ and $\mathbb{P}_{(\Pi)}^{(\mathcal{A})}$ is a genuine (non-negative scalar-valued) probability measure on $\mathbb{R}^{2K},$ and hence the integral on the right-hand-side of (6) is well-defined and takes values in $[0,+\infty].$ In view of (5), the loss function $\mathbb{E}_{(\Pi)}^{(\mathcal{A})}\left[c(X,Y)\right]$ is equal to
@@
In transport theory, a usual “finite energy condition” ensuring that $\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x,y)\mathrm{d}\pi(x,y)$ is finite for every coupling of $\mu$ and $\nu$ is that $c(x_{1},\dots,x_{K},y_{1},\dots,y_{K})\leq c_{X}(x_{1},\dots,x_{K})+c_{Y}(y_{1},\dots,y_{K})$ for some $c_{X}\in L^{1}(\mathbb{R}^{K},\mu)$ and $c_{Y}\in L^{1}(\mathbb{R}^{K},\nu)$ — see, e.g., [49, Remark 4.5]. Accordingly, here we impose the following finite energy condition. Let $\mu_{k}$ and $\nu_{k}$ be the laws of $A_{k}$ in the states $\rho$ and $\omega$ respectively, that is, $\mathrm{d}\mu_{k}(x_{k})=\mathrm{tr}_{\mathcal{H}}[\rho\,\mathrm{d}E_{k}(x_{k})]$ and $\mathrm{d}\nu_{k}(y_{k})=\mathrm{tr}_{\mathcal{H}}[\omega\,\mathrm{d}E_{k}(y_{k})]$ for every $k\in\{1,\dots,K\},$ and let $c(x,y)$ satisfy $0\leq c(x,y)\leq c_{X}(x_{1},\dots,x_{K})+c_{Y}(y_{1},\dots,y_{K})$ for some $c_{X}\in L^{1}(\mathbb{R}^{K},\mu_{1}\otimes\dots\otimes\mu_{K})$ and $c_{Y}\in L^{1}(\mathbb{R}^{K},\nu_{1}\otimes\dots\otimes\nu_{K}).$ Under this assumption, $\mathbb{E}_{(\Pi)}^{(\mathcal{A})}[c(X,Y)]$ is finite for any $\pi\in\mathcal{C}(\rho,\omega).$ Indeed, in this case
@@
$=\int_{\mathbb{R}^{K}}c_{X}(x_{1},\dots,x_{K})\prod_{k=1}^{K}\mathrm{tr}_{\mathcal{H}}[\rho\,\mathrm{d}E_{k}(x_{k})]+\int_{\mathbb{R}^{K}}c_{Y}(y_{1},\dots,y_{K})\prod_{k=1}^{K}\mathrm{tr}_{\mathcal{H}}[\omega\,\mathrm{d}E_{k}(y_{k})]$
$=\int_{\mathbb{R}^{K}}c_{X}\mathrm{d}(\mu_{1}\otimes\dots\otimes\mu_{K})+\int_{\mathbb{R}^{K}}c_{Y}\mathrm{d}(\nu_{1}\otimes\dots\otimes\nu_{K})<\infty.$
@@
[…]
where $\Pi$ runs over the set of all couplings of $\rho$ and $\omega$ which are states on $\mathcal{H}.$ As the cost operator $C_{c}^{(\mathcal{A})}$ acting on its domain contained in $\left(\mathcal{H}\otimes\mathcal{H}^{*}\right)^{\otimes K}$ is unbounded if so is $c$ on $\mathbb{R}^{2K},$ we emphasize that by the formal expression $\mathrm{tr}_{\left(\mathcal{H}\otimes\mathcal{H}^{*}\right)^{\otimes K}}\left[\Pi^{\otimes K}C_{c}^{(\mathcal{A})}\right]$ we mean $\lim_{n\to\infty}\mathrm{tr}_{\left(\mathcal{H}\otimes\mathcal{H}^{*}\right)^{\otimes K}}\left[\Pi^{\otimes K}\chi_{n}\left(C_{c}^{(\mathcal{A})}\right)\right],$ where $\chi_{n}:[0,\infty)\rightarrow[0,n]$ is the cut function defined by $\chi_{n}(x)=\min{\{x,n\}}.$ Note that this limit exists in $[0,+\infty]$ and coincides with $\mathbb{E}_{(\Pi)}^{(\mathcal{A})}[c(X,Y)]=\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x,y)\mathrm{d}\mathbb{P}_{(\Pi)}^{(\mathcal{A})}(x,y).$
@@
Then, according to the compactness result concerning quantum transport plans [16, Proposition 8], there is a subsequence $\left\{\Pi_{n_{j}}\right\}_{j\in\mathbb{N}}$ and an element $\widetilde{\Pi}\in\mathcal{C}(\rho,\omega)$ such that $\Pi_{n_{j}}\to\widetilde{\Pi}$ in trace norm. Consequently, $\mathbb{P}_{\left(\Pi_{n_{j}}\right)}^{(\mathcal{A})}$ converges to $\mathbb{P}_{\left(\widetilde{\Pi}\right)}^{(\mathcal{A})}$ set-wise. Indeed, for any Borel set $B\subset\mathbb{R}^{2K}$ the convergence
@@
follows from the fact that $\mathrm{tr}_{(\mathcal{H}\otimes\mathcal{H}^{*})^{\otimes K}}\left[\left|\left(\Pi_{n_{j}}\right)^{\otimes K}-\widetilde{\Pi}^{\otimes K}\right|\right]$ tends to 00 as $j$ tends to infinity. The last step will be the proof of the lower semi-continuity of the cost functional
@@
on the set of Borel probabilities on $\mathbb{R}^{K}\times\mathbb{R}^{K}$ with respect to the set-wise convergence. This is a standard argument (see, e.g., [49, Lemma 4.3]) which is based on the lower semi-continuity of $(x,y)\mapsto c(x,y)$ itself. As $c$ is non-negative and lower semi-continuous, it is the pointwise limit of a monotone increasing family of non-negative, continuous and bounded functions $\left\{c_{r}\right\}_{r\in\mathbb{N}}$ on $\mathbb{R}^{K}.$ By Beppo Levi’s monotone convergence theorem,
@@
$=\lim_{r\to\infty}\left(\lim_{j\to\infty}\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c_{r}(x,y)\mathrm{d}\mathbb{P}_{\left(\Pi_{n_{j}}\right)}^{(\mathcal{A})}(x,y)\right)\leq\liminf_{j\to\infty}\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x,y)\mathrm{d}\mathbb{P}_{\left(\Pi_{n_{j}}\right)}^{(\mathcal{A})}(x,y).$
@@
$=\left(\sum_{k=1}^{K}\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}\left|x_{k}-y_{k}\right|^{q}\mathrm{d}E_{1}(y_{1})\otimes\mathrm{d}E_{1}^{T}(x_{1})\otimes\dots\otimes\mathrm{d}E_{K}(y_{K})\otimes\mathrm{d}E_{K}^{T}(x_{K})\right)^{\frac{p}{q}}$
$=\left(\sum_{k=1}^{K}\left|\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}\left(x_{k}-y_{k}\right)\mathrm{d}E_{1}(y_{1})\otimes\mathrm{d}E_{1}^{T}(x_{1})\otimes\dots\otimes\mathrm{d}E_{K}(y_{K})\otimes\mathrm{d}E_{K}^{T}(x_{K})\right|^{q}\right)^{\frac{p}{q}}$
$=\left(\sum_{k=1}^{K}\left(\left|A_{k}\otimes I_{\mathcal{H}}^{T}-I_{\mathcal{H}}\otimes A_{k}^{T}\right|^{q}\right)^{(k)}\right)^{\frac{p}{q}}$ (10)
@@
$=\sum_{k=1}^{K}\mathrm{tr}_{\left(\mathcal{H}\otimes\mathcal{H}^{*}\right)}\left[\Pi\left|A_{k}\otimes I_{\mathcal{H}}^{T}-I_{\mathcal{H}}\otimes A_{k}^{T}\right|^{p}\right].$ (12)
@@
In the even more special case $p=q=2$ we get back the definition of the quadratic quantum Wasserstein distance introduced by De Palma and Trevisan in [16, Definition 8].
@@ 2.2. A quantum mechanical optimal transport problem without imposing independence
$=I_{L^{2}(\mathbb{R})}\otimes\dots\otimes I_{L^{2}(\mathbb{R})}\otimes\int_{\mathbb{R}}\lambda\mathrm{d}E(\lambda)\otimes I_{L^{2}(\mathbb{R})}\otimes\dots\otimes I_{L^{2}(\mathbb{R})}$
$=\int_{\mathbb{R}}\lambda I_{L^{2}(\mathbb{R})}\otimes\dots\otimes I_{L^{2}(\mathbb{R})}\otimes\mathrm{d}E(\lambda)\otimes I_{L^{2}(\mathbb{R})}\otimes\dots\otimes I_{L^{2}(\mathbb{R})}=\int_{\mathbb{R}}\lambda\mathrm{d}E(\lambda)^{(k)}.$
@@
$=\mathrm{tr}_{L^{2}(\mathbb{R}^{K})\otimes\left(L^{2}(\mathbb{R}^{K})\right)^{*}}\left[\Pi\left(\mathrm{d}E(y_{1})\otimes\mathrm{d}E(y_{2})\otimes\dots\otimes\mathrm{d}E(y_{K})\otimes\mathrm{d}E(x_{1})^{T}\otimes\mathrm{d}E(x_{2})^{T}\otimes\dots\otimes\mathrm{d}E(x_{K})^{T}\right)\right].$ (17)
@@
$=\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x_{1},\dots,x_{K},y_{1},\dots,y_{K})\mathrm{d}\mathbb{P}(Y_{1}=y_{1},Y_{2}=y_{2},\dots,Y_{K}=y_{K},X_{1}=x_{1},X_{2}=x_{2},\dots,X_{K}=x_{K})$
$=\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x_{1},\dots,x_{K},y_{1},\dots,y_{K})\mathrm{tr}_{L^{2}(\mathbb{R}^{K})\otimes\left(L^{2}(\mathbb{R}^{K})\right)^{*}}\left[\Pi\left(\mathrm{d}E(y_{1})\otimes\dots\otimes\mathrm{d}E(y_{K})\otimes\mathrm{d}E(x_{1})^{T}\otimes\dots\otimes\mathrm{d}E(x_{K})^{T}\right)\right]$
$=\mathrm{tr}_{L^{2}(\mathbb{R}^{K})\otimes\left(L^{2}(\mathbb{R}^{K})\right)^{*}}\left[\Pi\left(\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x_{1},\dots,x_{K},y_{1},\dots,y_{K})\mathrm{d}E(y_{1})\otimes\dots\otimes\mathrm{d}E(y_{K})\otimes\mathrm{d}E(x_{1})^{T}\otimes\dots\otimes\mathrm{d}E(x_{K})^{T}\right)\right].$
@@
minimize $\text{minimize }\mathrm{tr}_{L^{2}(\mathbb{R}^{K})\otimes\left(L^{2}(\mathbb{R}^{K})\right)^{*}}\left[\Pi C_{c}\right]\text{ over all }\Pi\in\mathcal{C}(\rho,\omega).$ (20)
@@
$=\left(\sum_{k=1}^{K}\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}\left|x_{k}-y_{k}\right|^{q}\mathrm{d}E(y_{1})\otimes\dots\otimes\mathrm{d}E(y_{K})\otimes\mathrm{d}E(x_{1})^{T}\otimes\dots\otimes\mathrm{d}E(x_{K})^{T}\right)^{\frac{p}{q}}$
$=\left(\sum_{k=1}^{K}\left|\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}\left(x_{k}-y_{k}\right)\mathrm{d}E(y_{1})\otimes\dots\otimes\mathrm{d}E(y_{K})\otimes\mathrm{d}E(x_{1})^{T}\otimes\dots\otimes\mathrm{d}E(x_{K})^{T}\right|^{q}\right)^{\frac{p}{q}}$
$=\left(\sum_{k=1}^{K}\left|Q^{(k)}\otimes I_{L^{2}(\mathbb{R}^{K})}^{T}-I_{L^{2}(\mathbb{R}^{K})}\otimes\left(Q^{(k)}\right)^{T}\right|^{q}\right)^{\frac{p}{q}}.$

### v3

@@ Abstract.
Bunth was supported by the Momentum Program of the Hungarian Academy of Sciences (grant no. LP2021-15/2021); Pitrik was supported by the “Frontline” Research Excellence Programme of the Hungarian National Research, Development and Innovation Office - NKFIH (grant no. KKP133827) and by the Momentum Program of the Hungarian Academy of Sciences (grant no. LP2021-15/2021); Titkos was supported by the Hungarian National Research, Development and Innovation Office - NKFIH (grant no. K115383) and by the Momentum Program of the Hungarian Academy of Sciences (grant no. LP2021-15/2021).; Virosztek was supported by the Momentum program of the Hungarian Academy of Sciences under grant agreement no. LP2021-15/2021, by the Hungarian National Research, Development and Innovation Office (NKFIH) under grant agreement no. Excellence_151232, and partially supported by the ERC Synergy Grant No. 810115.
@@ 1.1. Motivation and main result
[…]
@@ 1.2. Basic notions, notation
where $\pi$ runs over all possible couplings of $\mu$ and $\nu.$ We say that $\pi\in\mathrm{Prob}(\mathbb{R}^{K}\times\mathbb{R}^{K})$ is a coupling of $\mu$ and $\nu$ (in notation: $\pi\in\mathcal{C}(\mu,\nu)$) if the marginals of $\pi$ are $\mu$ and $\nu,$ that is, $\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}f(x)\mathrm{d}\pi(x,y)=\int_{\mathbb{R}^{K}}f(x)\mathrm{d}\mu(x)$ and $\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}g(y)\mathrm{d}\pi(x,y)=\int_{\mathbb{R}^{K}}g(y)\mathrm{d}\nu(y)$ for all $f,g\in C_{b}(\mathbb{R}^{K}).$ A consequence of the tightness (that is, sequential compactness in the weak topology) of $\mathcal{C}(\mu,\nu)$ and the lower-semicontinuity of $c$ is that there is a coupling (in other words: transport plan) $\pi_{0}\in\mathcal{C}(\mu,\nu)$ that minimizes (1), see, e.g., [50, Thm. 4.1.].
@@ 2.1. A quantum mechanical optimal transport problem generalizing [17]
$\prod_{k=1}^{K}\mathrm{tr}_{\mathcal{H}\otimes\mathcal{H}^{*}}\left[\Pi\left(\mathrm{d}E_{k}(y_{k})\otimes\mathrm{d}E_{k}^{T}(x_{k})\right)\right]$
$\mathrm{tr}_{\left(\mathcal{H}\otimes\mathcal{H}^{*}\right)^{\otimes K}}\left[\Pi^{\otimes K}\left(\mathrm{d}E_{1}(y_{1})\otimes\mathrm{d}E_{1}^{T}(x_{1})\otimes\dots\otimes\mathrm{d}E_{K}(y_{K})\otimes\mathrm{d}E_{K}^{T}(x_{K})\right)\right].$ (5)
@@
Note that $c$ is a non-negative Borel measurable function on $\mathbb{R}^{2K},$ and $\mathbb{P}_{(\Pi)}^{(\mathcal{A})}$ is a genuine (non-negative scalar-valued) probability measure on $\mathbb{R}^{2K},$ and hence the integral on the right-hand-side of (6) is well-defined and takes values in $0,+\infty].$ In view of (5), the loss function $\mathbb{E}_{(\Pi)}^{(\mathcal{A})}\left[c(X,Y)\right]$ is equal to
@@
In transport theory, a usual “finite energy condition” ensuring that $\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x,y)\mathrm{d}\pi(x,y)$ is finite for every coupling of $\mu$ and $\nu$ is that $c(x_{1},\dots,x_{K},y_{1},\dots,y_{K})\leq c_{X}(x_{1},\dots,x_{K})+c_{Y}(y_{1},\dots,y_{K})$ for some $c_{X}\in L^{1}(\mathbb{R}^{K},\mu)$ and $c_{Y}\in L^{1}(\mathbb{R}^{K},\nu)$ — see, e.g., [50, Remark 4.5]. Accordingly, here we impose the following finite energy condition. Let $\mu_{k}$ and $\nu_{k}$ be the laws of $A_{k}$ in the states $\rho$ and $\omega$ respectively, that is, $\mathrm{d}\mu_{k}(x_{k})=\mathrm{tr}_{\mathcal{H}}[\rho\,\mathrm{d}E_{k}(x_{k})]$ and $\mathrm{d}\nu_{k}(y_{k})=\mathrm{tr}_{\mathcal{H}}[\omega\,\mathrm{d}E_{k}(y_{k})]$ for every $k\in\{1,\dots,K\},$ and let $c(x,y)$ satisfy $0\leq c(x,y)\leq c_{X}(x_{1},\dots,x_{K})+c_{Y}(y_{1},\dots,y_{K})$ for some $c_{X}\in L^{1}(\mathbb{R}^{K},\mu_{1}\otimes\dots\otimes\mu_{K})$ and $c_{Y}\in L^{1}(\mathbb{R}^{K},\nu_{1}\otimes\dots\otimes\nu_{K}).$ Under this assumption, $\mathbb{E}_{(\Pi)}^{(\mathcal{A})}[c(X,Y)]$ is finite for any $\pi\in\mathcal{C}(\rho,\omega).$ Indeed, in this case
@@
$\int_{\mathbb{R}^{K}}c_{X}(x_{1},\dots,x_{K})\prod_{k=1}^{K}\mathrm{tr}_{\mathcal{H}}[\rho\,\mathrm{d}E_{k}(x_{k})]+\int_{\mathbb{R}^{K}}c_{Y}(y_{1},\dots,y_{K})\prod_{k=1}^{K}\mathrm{tr}_{\mathcal{H}}[\omega\,\mathrm{d}E_{k}(y_{k})]$
$\int_{\mathbb{R}^{K}}c_{X}\mathrm{d}(\mu_{1}\otimes\dots\otimes\mu_{K})+\int_{\mathbb{R}^{K}}c_{Y}\mathrm{d}(\nu_{1}\otimes\dots\otimes\nu_{K})<\infty.$
@@
 $\text{ minimize }\Pi\mapsto\mathrm{tr}_{\left(\mathcal{H}\otimes\mathcal{H}^{*}\right)^{\otimes K}}\left[\Pi^{\otimes K}C_{c}^{(\mathcal{A})}\right]$ (8)
where $\Pi$ runs over the set of all couplings of $\rho$ and $\omega$ which are states on $\mathcal{H}.$ As the cost operator $C_{c}^{(\mathcal{A})}$ acting on its domain contained in $\left(\mathcal{H}\otimes\mathcal{H}^{*}\right)^{\otimes K}$ is unbounded if so is $c$ on $\mathbb{R}^{2K},$ we emphasize that by the formal expression $\mathrm{tr}_{\left(\mathcal{H}\otimes\mathcal{H}^{*}\right)^{\otimes K}}\left[\Pi^{\otimes K}C_{c}^{(\mathcal{A})}\right]$ we mean $\lim_{n\to\infty}\mathrm{tr}_{\left(\mathcal{H}\otimes\mathcal{H}^{*}\right)^{\otimes K}}\left[\Pi^{\otimes K}\chi_{n}\left(C_{c}^{(\mathcal{A})}\right)\right],$ where $\chi_{n}:[0,\infty)\rightarrow[0,n]$ is the cut function defined by $\chi_{n}(x)=\min{\{x,n\}}.$ Note that this limit exists in $0,+\infty]$ and coincides with $\mathbb{E}_{(\Pi)}^{(\mathcal{A})}[c(X,Y)]=\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x,y)\mathrm{d}\mathbb{P}_{(\Pi)}^{(\mathcal{A})}(x,y).$
@@
Then, according to the compactness result concerning quantum transport plans [17, Proposition 8], there is a subsequence $\left\{\Pi_{n_{j}}\right\}_{j\in\mathbb{N}}$ and an element $\widetilde{\Pi}\in\mathcal{C}(\rho,\omega)$ such that $\Pi_{n_{j}}\to\widetilde{\Pi}$ in trace norm. Consequently, $\mathbb{P}_{\left(\Pi_{n_{j}}\right)}^{(\mathcal{A})}$ converges to $\mathbb{P}_{\left(\widetilde{\Pi}\right)}^{(\mathcal{A})}$ set-wise. Indeed, for any Borel set $B\subset\mathbb{R}^{2K}$ the convergence
@@
follows from the fact that $\mathrm{tr}_{(\mathcal{H}\otimes\mathcal{H}^{*})^{\otimes K}}\left[\left|\left(\Pi_{n_{j}}\right)^{\otimes K}-\widetilde{\Pi}^{\otimes K}\right|\right]$ tends to 0 as $j$ tends to infinity. The last step will be the proof of the lower semi-continuity of the cost functional
@@
on the set of Borel probabilities on $\mathbb{R}^{K}\times\mathbb{R}^{K}$ with respect to the set-wise convergence. This is a standard argument (see, e.g., [50, Lemma 4.3]) which is based on the lower semi-continuity of $(x,y)\mapsto c(x,y)$ itself. As $c$ is non-negative and lower semi-continuous, it is the pointwise limit of a monotone increasing family of non-negative, continuous and bounded functions $\left\{c_{r}\right\}_{r\in\mathbb{N}}$ on $\mathbb{R}^{K}.$ By Beppo Levi’s monotone convergence theorem,
@@
$\lim_{r\to\infty}\left(\lim_{j\to\infty}\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c_{r}(x,y)\mathrm{d}\mathbb{P}_{\left(\Pi_{n_{j}}\right)}^{(\mathcal{A})}(x,y)\right)\leq\liminf_{j\to\infty}\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x,y)\mathrm{d}\mathbb{P}_{\left(\Pi_{n_{j}}\right)}^{(\mathcal{A})}(x,y).$
@@
$\left(\sum_{k=1}^{K}\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}\left|x_{k}-y_{k}\right|^{q}\mathrm{d}E_{1}(y_{1})\otimes\mathrm{d}E_{1}^{T}(x_{1})\otimes\dots\otimes\mathrm{d}E_{K}(y_{K})\otimes\mathrm{d}E_{K}^{T}(x_{K})\right)^{\frac{p}{q}}$
$\left(\sum_{k=1}^{K}\left|\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}\left(x_{k}-y_{k}\right)\mathrm{d}E_{1}(y_{1})\otimes\mathrm{d}E_{1}^{T}(x_{1})\otimes\dots\otimes\mathrm{d}E_{K}(y_{K})\otimes\mathrm{d}E_{K}^{T}(x_{K})\right|^{q}\right)^{\frac{p}{q}}$
$\left(\sum_{k=1}^{K}\left(\left|A_{k}\otimes I_{\mathcal{H}}^{T}-I_{\mathcal{H}}\otimes A_{k}^{T}\right|^{q}\right)^{(k)}\right)^{\frac{p}{q}}$ (10)
@@
$\sum_{k=1}^{K}\mathrm{tr}_{\left(\mathcal{H}\otimes\mathcal{H}^{*}\right)}\left[\Pi\left|A_{k}\otimes I_{\mathcal{H}}^{T}-I_{\mathcal{H}}\otimes A_{k}^{T}\right|^{p}\right].$ (12)
@@
In the even more special case $p=q=2$ we get back the definition of the quadratic quantum Wasserstein distance introduced by De Palma and Trevisan in [17, Definition 8].
@@ 2.2. A quantum mechanical optimal transport problem without imposing independence
$I_{L^{2}(\mathbb{R})}\otimes\dots\otimes I_{L^{2}(\mathbb{R})}\otimes\int_{\mathbb{R}}\lambda\mathrm{d}E(\lambda)\otimes I_{L^{2}(\mathbb{R})}\otimes\dots\otimes I_{L^{2}(\mathbb{R})}$
$\int_{\mathbb{R}}\lambda I_{L^{2}(\mathbb{R})}\otimes\dots\otimes I_{L^{2}(\mathbb{R})}\otimes\mathrm{d}E(\lambda)\otimes I_{L^{2}(\mathbb{R})}\otimes\dots\otimes I_{L^{2}(\mathbb{R})}=\int_{\mathbb{R}}\lambda\mathrm{d}E(\lambda)^{(k)}.$
@@
$\mathrm{tr}_{L^{2}(\mathbb{R}^{K})\otimes\left(L^{2}(\mathbb{R}^{K})\right)^{*}}\left[\Pi\left(\mathrm{d}E(y_{1})\otimes\mathrm{d}E(y_{2})\otimes\dots\otimes\mathrm{d}E(y_{K})\otimes\mathrm{d}E(x_{1})^{T}\otimes\mathrm{d}E(x_{2})^{T}\otimes\dots\otimes\mathrm{d}E(x_{K})^{T}\right)\right].$ (17)
@@
$\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x_{1},\dots,x_{K},y_{1},\dots,y_{K})\mathrm{d}\mathbb{P}(Y_{1}=y_{1},Y_{2}=y_{2},\dots,Y_{K}=y_{K},X_{1}=x_{1},X_{2}=x_{2},\dots,X_{K}=x_{K})$
$\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x_{1},\dots,x_{K},y_{1},\dots,y_{K})\mathrm{tr}_{L^{2}(\mathbb{R}^{K})\otimes\left(L^{2}(\mathbb{R}^{K})\right)^{*}}\left[\Pi\left(\mathrm{d}E(y_{1})\otimes\dots\otimes\mathrm{d}E(y_{K})\otimes\mathrm{d}E(x_{1})^{T}\otimes\dots\otimes\mathrm{d}E(x_{K})^{T}\right)\right]$
$\mathrm{tr}_{L^{2}(\mathbb{R}^{K})\otimes\left(L^{2}(\mathbb{R}^{K})\right)^{*}}\left[\Pi\left(\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}c(x_{1},\dots,x_{K},y_{1},\dots,y_{K})\mathrm{d}E(y_{1})\otimes\dots\otimes\mathrm{d}E(y_{K})\otimes\mathrm{d}E(x_{1})^{T}\otimes\dots\otimes\mathrm{d}E(x_{K})^{T}\right)\right].$
@@
$\text{minimize }\mathrm{tr}_{L^{2}(\mathbb{R}^{K})\otimes\left(L^{2}(\mathbb{R}^{K})\right)^{*}}\left[\Pi C_{c}\right]\text{ over all }\Pi\in\mathcal{C}(\rho,\omega).$ (20)
@@
$\left(\sum_{k=1}^{K}\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}\left|x_{k}-y_{k}\right|^{q}\mathrm{d}E(y_{1})\otimes\dots\otimes\mathrm{d}E(y_{K})\otimes\mathrm{d}E(x_{1})^{T}\otimes\dots\otimes\mathrm{d}E(x_{K})^{T}\right)^{\frac{p}{q}}$
$\left(\sum_{k=1}^{K}\left|\iint_{\mathbb{R}^{K}\times\mathbb{R}^{K}}\left(x_{k}-y_{k}\right)\mathrm{d}E(y_{1})\otimes\dots\otimes\mathrm{d}E(y_{K})\otimes\mathrm{d}E(x_{1})^{T}\otimes\dots\otimes\mathrm{d}E(x_{K})^{T}\right|^{q}\right)^{\frac{p}{q}}$
$\left(\sum_{k=1}^{K}\left|Q^{(k)}\otimes I_{L^{2}(\mathbb{R}^{K})}^{T}-I_{L^{2}(\mathbb{R}^{K})}\otimes\left(Q^{(k)}\right)^{T}\right|^{q}\right)^{\frac{p}{q}}.$